OpenStudy (anonymous):

Find the force applied by car brakes knowing that the mass of the car is 3000 kg, velocity 20 m/s and that the brakes applied 55 meters before the stoplight

OpenStudy (anonymous):

use the work-energy relation \[ \frac{1}{2} m v^2 = F s \] so we get \[ F = \frac{1}{2}\frac{mv^2}{s} \]